A new electric saw for cutting small pieces of lumber in a furniture manufacturing plant has a cost
basis of Rs. 5,000 and 8 year depreciable life. The estimated salvage value of the saw is zero at the
end of 4 th year cumulative depreciation charge of 7 th year and book value of 6 th year using double
declining balance method.
1
Expert's answer
2020-05-26T11:29:32-0400

Let's calculate the book value (BV) and depreciation (D) during all 8 years:

Year 1: BV = 5,000, D = 5,000×2/8 = 1,250.

Year 2: BV = 4,750, D = 1,187.5,

Year 3: BV = 3,562.5, D = 890.6,

Year 4: BV = 2,671.9, D = 668,

Year 5: BV = 2,003.9, D = 501,

Year 6: BV = 1,502.9, D = 375.7,

Year 7: BV = 1,127.2, D = 281.8,

The cumulative depreciation charge of 7 th year is: 1,187.5 + 890.6 + 668 + 501 + 375.7 + 281.8 = 3904.6,

and book value of 6 th year is 1,502.9 using double declining balance method.